Another problem I just couldn't figure out and I'm still working on it . A product that cost 2,000 increase by 5% in June. The new June price of the product increased by 5% on December 1st. How many dollars did the product increase by Dec 1. I don't have a clue! Can you please help? Sorry but, I'm still having trouble

Answer by ankor@dixie-net.com(22740)

A product that cost 2,000 increase by 5% in June. The new June price of the product increased by 5% on December 1st. How many dollars did the product increase by Dec 1.
:
Find the cost of the item after the 1st 5% increase (june)
1.05 * 2000 = 2100
:
Find the cost of the item after the 2nd 5% increase (dec)
1.05 * 2100 = $2205
:
What was the total increase?
2205 - 2000 = $205